what is the density of (CCl4) vapor at 0.881721 atm and 125 celcius? (in g/L)
Answer #1

P = 0.881721 atm
T= 125.0 oC
= (125.0+273) K
= 398 K

Molar mass of CCl4,
MM = 1*MM(C) + 4*MM(Cl)
= 1*12.01 + 4*35.45
= 153.81 g/mol

Lets derive the equation to be used
use:
p*V=n*R*T
p*V=(mass/molar mass)*R*T
p*molar mass=(mass/V)*R*T
p*molar mass=density*R*T

Put Values:
0.881721 atm *153.81 g/mol = density * 0.08206 atm.L/mol.K *398.0 K
density = 4.1524 g/L
Answer: 4.15 g/L